A rectangle has a diagonal of 5 meters and perimeter of 14 meters. Find the length and width of the rectangle.
Answer by jim_thompson5910(35256)

P = 2L+2W

14 = 2L+2W

7 = L+W

L = 7-W

-----------------
L^2+W^2 = 5^2

L^2+W^2 = 25

(7-W)^2+W^2 = 25

49-14W+W^2+W^2 - 25 = 0

2W^2-14W+24 = 0

W^2-7W+12=0

(W-4)(W-3) = 0

W-4=0 or W-3=0

W = 4 or W = 3
----------------
So the width is 3 meters and the length is 4 meters.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------